Please solve it by hand or typing, don't use EXCEL.

1. The following data was downloaded from Professor Kenneth Frenchs website. The table shows the annual return of the U.S. stock market including NYSE, NASDAQ and Amex over the 10 years. (15pts) Year 2000 2001 2002 2003 2004 2005 2006 2007 2008 2009 Return (%) -16.97-15.12-22.47 32.12 11.82 4.34 11.4 263-39 96 3158 1) What would be your expected return of the stock market in 2010 based on the table? (4pts) 2) What would be your expected standard deviation of return in 2010? (4pts) 3) The average T-bill rate was 2.77% annually over the last 10 years, what would be the Sharpe ratio of the U.S. stock market? (3pts) 4) What would be the 10% value at risk if the return follows a normal distribution? (2pts) 5) Explain the meaning of your answer to question d. (2pts)
Answer #1

1) Expected Return = Compounded Annual Growth Rate (CAGR) over the past years

Considering in 1999, the index value was 100, the index value in 2009 is

100*(1-0.1697)*(1-0.1512)*(1-0.2248)*(1+0.3212)*(1+0.1182)*(1+0.0434)*(1+0.114)*(1+0.0263)*(1-0.3996)*(1+0.3158) = 76.07

CAGR = (76.07/100)^(1/10) - 1 = -2.70%

Therefore, expected return for 2010 = -2.70%

2 Std. Dev of historical returns = 23.43%

Therefore, expected standard deviation = 23.43%

3. Sharpe Ratio = (Market Return - Treasury Return) / Standard Deviation = (-0.0270 - 0.0277)/(0.2343) = -0.15

4. 10% VAR for normal distribution =

At significance level of 10%, z=1.28

(x-76.07*(1-0.027))/0.02343 = 1.28

x= 74.32

5. Therefore, there is a 10% probability that index value will fall to 74.32 or below.